# STRIDE: Single-video based Temporally Continuous Occlusion Robust 3D Pose Estimation

### v2

@@ Abstract
The capability to accurately estimate 3D human poses is crucial for diverse fields such as action recognition, gait recognition, and virtual/augmented reality. However, a persistent and significant challenge within this field is the accurate prediction of human poses under conditions of severe occlusion. Traditional image-based estimators struggle with heavy occlusions due to a lack of temporal context, resulting in inconsistent predictions. While video-based models benefit from processing temporal data, they encounter limitations when faced with prolonged occlusions that extend over multiple frames. This challenge arises because these models struggle to generalize beyond their training datasets, and the variety of occlusions is hard to capture in the training data. Addressing these challenges, we propose STRIDE (Single-video based TempoRally contInuous occlusion Robust 3D Pose Estimation), a novel Test-Time Training (TTT) approach to fit a human motion prior for each video. This approach specifically handles occlusions that were not encountered during the model’s training. By employing STRIDE, we can refine a sequence of noisy initial pose estimates into accurate, temporally coherent poses during test time, effectively overcoming the limitations of prior methods. Our framework demonstrates flexibility by being model-agnostic, allowing us to use any off-the-shelf 3D pose estimation method for improving robustness and temporal consistency. We validate STRIDE’s efficacy through comprehensive experiments on challenging datasets like Occluded Human3.6M, Human3.6M, and OCMotion, where it not only outperforms existing single-image and video-based pose estimation models but also showcases superior handling of substantial occlusions, achieving fast, robust, accurate, and temporally consistent 3D pose estimates.
00footnotetext: * Equal Contribution
Figure 1: Effect of occlusions on pose estimation. Image-based 3D pose estimators [3] often struggle with heavy occlusions, as illustrated in this figure. Without temporal context, predictions on highly obscured frames are inconsistent with prior poses, like the erroneous pose in the third frame. Notably, even state-of-the-art video approaches [39] fail on prolonged full occlusions spanning multiple frames, as in frames 4-5. This highlights another critical limitation - models are brittle when deployed outside their training distributions. Without training examples of such long-duration occlusions, models fail to extrapolate reasonable poses. Our work addresses this through test-time training of a human motion prior. By fine-tuning on each new video, we tailor this parametric prior to handling sequence-specific occlusion patterns not observed during training. Given an initial noisy estimate, our approach refines the pose sequence into an accurate, temporally coherent output, as shown in the final row.
@@ 1 Introduction
Accurate 3D pose estimation [48] is an important problem in computer vision with a variety of real-world applications, including but not limited to action recognition [19], virtual and augmented reality [1], and gait recognition [51, 10]. While the performance of 3D pose estimation algorithms has improved rapidly in recent years, the majority of these are image-based [37, 33, 35], estimating the pose from a single image. Consequently, these approaches still face inherent challenges in handling occluded subjects due to the limited visual information contained in individual images. To address these issues, recent efforts have explored video-based pose estimation algorithms [43, 36], leveraging temporal continuity across frames to resolve pose ambiguities from missing visual evidence.
@@
Recent TTT approaches for 3D pose estimation [30, 29, 12, 11] fine-tune the model using 2D cues like keypoints or silhouettes extracted from the test images. However, this reliance on 2D cues has inherent limitations. Firstly, the 2D projection of 3D poses is ambiguous, as many plausible 3D configurations can map to the same 2D keypoints. Secondly, 2D pose estimators themselves are susceptible to errors on unseen data distributions [35, 17]. Thus, fine-tuning on potentially imperfect and ambiguous 2D poses can incorrectly update the model, leading to degraded 3D predictions.
@@
To overcome the limitations of existing methods, we propose STRIDE (Single-video based TempoRally contInuous occlusion Robust 3D Pose Estimation), a novel test-time training framework for 3D pose estimation under occlusion. The key component of our approach is a parametric motion prior that is capable of modeling the dynamics of natural human motions and poses. This motion prior is pre-trained using a BERT-style [9, 52] approach on 3D pose sequences, learning to reconstruct temporally coherent poses when given a series of noisy estimates as input. At test time, given a sequence of noisy 3D poses from any existing pose estimation algorithm, STRIDE leverages this pre-trained prior to produce a clean sequence by fine-tuning it on each new video. We use 3D kinematic losses for motion smoothing via adapting the model to the video-specific motion patterns. By leveraging the motion prior’s inherent knowledge of natural human movement during test-time training, STRIDE avoids ambiguities of 2D pose information faced by existing approaches. An overview of our approach is shown in Fig. 2.
A key advantage of our algorithm is that it can work alongside any off-the-shelf pose estimator to improve temporal consistency, providing model-agnostic pose enhancements. This allows STRIDE to not only surpass image-based pose estimators that lack contextual cues to resolve occlusions, but also outperform video-based methods. Notably, STRIDE can handle situations with up to 100% occlusion of the human body over many consecutive frames. In comparison to existing test-time video based pose estimation method [36, 30], our approach is up to 2 times faster than previous state-of-the-art method [30] and operates without accessing any labeled training data during inference time, making it privacy [38] and storage-friendly. Contributions. In summary, we make the following key contributions:
@@
    We propose a novel test-time training method, STRIDE, for achieving temporally continuous 3D pose estimation under occlusion. This approach achieves robust performance without relying on labeled training data.
@@
    A motion prior model that refines noisy 3D pose sequences into smooth and continuous predictions.
@@
    A model-agnostic framework that can refine poses from any off-the-shelf estimator, highlighting efficiency and generalizability.
  4. 4.
    State-of-the-art results on challenging benchmarks including Occluded Human3.6M, Human3.6M [15], and OCMotion [14]. We demonstrate enhanced occlusion robustness and temporal consistency.
@@ 2 Related Works
Monocular 3D pose estimation. Monocular 3D pose estimation is a fundamental and challenging problem in computer vision which involves the localisation of 3D spatial pose coordinates from just a single image. The problem is inherently complex due to the diversity of body shapes, clothing, self-occlusions, etc. Despite these bottlenecks, recent deep learning-based methods have shown impressive performance on challenging academic datasets [3, 48]. [27] proposed the first CNN-based approach to regress 3D joints from a single image in an end-to-end fashion. Since then, numerous works [37, 33] have improved upon these ideas by using additional information such as multi-view constraints and depth information. Recent works [41] employ kinematic constraints for improved pose estimation, while [42] uses anatomical constraints and data augmentation for obtaining state-of-the-art results on academic datasets. However, it is worth noting that these methods are trained under supervised settings and often fail to generalize under distribution shifts. Owing to these weaknesses, [21, 22] proposed self-supervised algorithms for 3D human pose estimation. Although these works perform well in single image-based settings, they failed to generalise under occlusions and also lack temporal continuity when extended to video-based settings.
2D-3D human pose lifting. Modern 3D human pose estimation encounters significant challenges in generalization due to limited labeled data for real-world applications. [26] addressed this issue by breaking down the problem into 2D pose estimation and 2D to 3D lifting. Subsequently, [4] improved on this by including self-supervised geometric regularization, by synthetic data usage [53], spatio-temporal transformers [49], and frequency domain analysis [47]. [52] achieved state-of-the-art results by modelling motion priors from a sequence of 2D poses. Although these works perform well up to a certain degree, they suffer from two problems: 1) depth ambiguity of 2D human poses, 2) inaccurate 3D human poses if the initial 2D human poses are noisy. In contrast, we focus on 3D pose estimation in a video-based setting and does not involve any 2D-3D pose lifting.
@@
Video-based 3D pose estimation. Video based 3D human pose estimation have demonstrated impressive performance gains on challenging datasets. Work proposed in [50] performs direct regression to 3D human poses by employing consistency between 3D joints and 2D keypoints. [34] utilized temporal convolutions for 3D human pose estimation in videos. Works like [2] exploited SMPL pose and shape parameters from videos and used it for fine tuning HMR for improved human pose estimation in the wild. Further, [45] proposed a mixed spatio-temporal approach for 3D human pose regression which alternated between spatial consistency and temporal consistency. A recent method, HuMoR [36] performed a weighted regularization using predicted contact probabilities to maintain consistency among joint positions and joint heights across frames. The current state-of-the-art method CycleAdapt[30] handled the domain shift [20] between training and testing phases in 3D human mesh reconstruction by cyclically adapting a human mesh reconstruction network (HMRNet [16]) and a human motion denoising network (MDNet [30]) during test time. Despite the success of the above methods in maintaining temporal consistency, they are extremely slow due to an external optimization step and do not generalise well under distribution shifts. Severe occlusions often degrade the performance of these methods due to missing poses. Our work emphasizes these shortcomings and brings temporal continuity under severe occlusions by leveraging a motion-prior model that seamlessly handles missing poses.
@@
3D pose estimation under occlusion. Handling occlusions is a challenging problem, especially in video-based 3D pose estimation settings. [6] performed data augmentation using occlusion labels for 3D data using a novel Cylinder Man Model. Current methods solve this problem by refining the 3D poses to maintain temporal consistency. Recent methods like GLAMR [43] performed human mesh recovery in the global coordinate system from extracted motions in the local coordinate system and performed motion infilling for missing poses based on visible motions. SmoothNet [44] uses a temporal refinement network that takes poses from existing single image based pose estimation methods for alleviating motion jitters. Although these methods handle minor occlusions that infrequently occur in the scene, they do not perform well under heavy occlusions. To tackle this, our work adapts a motion prior from a sequence of noisy 3D poses to better predict the missing poses and maintain temporal consistency.
@@ 3 Method
Figure 3: The presented figure illustrates the pipeline for our temporally continuous pose estimation, STRIDE. Initially, we pre-train a motion prior model, denoted as $\mathcal{M}$, using a diverse set of 3D pose data sourced from various public datasets. The primary objective of this motion prior model is to generate a sequence of poses that exhibit temporal continuity when provided with a sequence of initially noisy poses. Moving into the single video training stage, we acquire a sequence of noisy poses using a 3D pose estimation model, $\mathcal{P}$. The weights of $\mathcal{P}$ are held constant during this phase. Subsequently, we pass this noisy pose sequence through the motion prior model $\mathcal{M}$ and retrain it using various supervised losses, as outlined in Eq. 5. The end result of this training process is a model capable of producing temporally continuous 3D poses for that specific video.
We address the problem of extracting temporally continuous 3D pose estimates from a monocular video that may contain heavy occlusions. Given an off-the-shelf monocular 3D pose estimator $\mathcal{P}$ (either image or video-based) that produces temporally inconsistent poses due to occlusions or domain gaps, our goal is to output clean, temporally coherent 3D pose sequences that better match natural human motion dynamics. To achieve this, we propose a two-stage approach, illustrated in Fig. 3:
@@
    Learning a motion prior: We first pre-train a self-attention-based motion prior model $\mathcal{M}$ on labeled 3D pose datasets in a BERT-style manner [9, 52]. During pre-training, we synthetically corrupt the 3D joint inputs with noise to simulate occlusions and other errors. $\mathcal{M}$ is then trained to denoise these inputs and reconstruct a sequence of temporally coherent 3D poses. This allows $\mathcal{M}$ to learn strong general priors of natural human motion dynamics.
@@
    Test-time alignment: For a given test video, we obtain potentially noisy per-frame poses using $\mathcal{P}$ [3] and fine-tune the motion prior model $\mathcal{M}$ in an unsupervised manner to align it to the specific motion exhibited in the video. This adaptation step allows us to obtain temporally continuous pose estimates for the given video.
In Section 3.1, we describe the architecture of the motion prior model $\mathcal{M}$. Next, in Section 3.2, we detail the masked sequence modelling approach used for pre-training the motion prior $\mathcal{M}$ on synthetically corrupted pose sequences. Finally, in Section 3.3, we introduce the self-supervised losses used for fine-tuning $\mathcal{M}$ at test time on each video.
@@ 3.1 Network Architecture
We base our motion prior model $\mathcal{M}$ on the DSTFormer architecture [52], originally proposed for lifting 2D poses to 3D. Here, we modify and adapt DSTFormer for the sequence-to-sequence task of denoising and smoothing noisy 3D pose sequence inputs. Specifically, the motion prior $\mathcal{M}$ takes in a sequence of 3D body poses represented as $\mathbf{X}\in\mathbb{R}^{T\times J\times 3}$, where $T$ is the number of frames, $J$ is the number of joints, and each pose consists of $J\times 3$ coordinate values. It then denoises the input sequence to produce refined temporally coherent 3D poses $\bar{\mathbf{X}}\in\mathbb{R}^{T\times J\times 3}$. $\mathcal{M}$ contains two key components: 1) a spatial block to capture the orientation of joints, and 2) a temporal block to model the temporal dynamics of a joint. The spatial block refines poses in each frame, while the temporal block smooths the transitions between frames. We describe these components below:
Spatial block. This block utilizes Spatial Multi-Head Self-Attention (S-MHSA) to model relationships among joints within each pose in the input sequence. Mathematically, the S-MHSA operation is defined as:
[…]
Here, $\mathbf{Q}_{\text{S}}^{i},\mathbf{K}_{\text{S}}^{i},\mathbf{V}_{\text{S}}^{i}$ denote the query, key, and value projections for the $i^{th}$ attention head, $d_{k}$ is the key dimension, and $\mathbf{W}_{\text{S}}^{P}$ is the projection parameter matrix. We apply S-MHSA to features of different time steps in parallel. The output undergoes further processing, including residual connection and layer normalization (LayerNorm), followed by a multi-layer perceptron (MLP).
Temporal block. This block utilizes Temporal Multi-Head Self-Attention (T-MHSA) to model the relationships between poses across time steps, thereby enabling the smoothing of the pose trajectories over the sequence. It operates similarly to S-MHSA but is applied to per-joint temporal features parallelized over the spatial dimension:
[…]
By attending to temporal relationships, T-MHSA produces smooth pose transitions over time.
Dual-Stream Spatio-temporal Transformer. We then use the dual-stream architecture which employs spatial and temporal Multi-Head Self-Attention mechanisms. These mechanisms capture intra-frame and inter-frame body joint interactions, necessitating careful consideration of three key assumptions: both streams model comprehensive spatio-temporal contexts, each stream specializes in distinct spatio-temporal aspects, and the fusion dynamically balances weights based on input characteristics.

### v4

# STRIDE: Single-video based Temporally Continuous Occlusion-Robust 3D Pose Estimation

@@ Abstract
Accurately estimating 3D human poses is crucial for fields like action recognition, gait recognition, and virtual/augmented reality. However, predicting human poses under severe occlusion remains a persistent and significant challenge. Existing image-based estimators struggle with heavy occlusions due to a lack of temporal context, resulting in inconsistent predictions, while video-based models, despite benefiting from temporal data, face limitations with prolonged occlusions over multiple frames. Additionally, existing algorithms often struggle to generalize unseen videos. Addressing these challenges, we propose STRIDE (Single-video based TempoRally contInuous Occlusion-Robust 3D Pose Estimation), a novel Test-Time Training (TTT) approach to fit a human motion prior for estimating 3D human poses for each video. Our proposed approach handles occlusions not encountered during the model’s training by refining a sequence of noisy initial pose estimates into accurate, temporally coherent poses at test time, effectively overcoming the limitations of existing methods. Our flexible, model-agnostic framework allows us to use any off-the-shelf 3D pose estimation method to improve robustness and temporal consistency. We validate STRIDE’s efficacy through comprehensive experiments on multiple challenging datasets where it not only outperforms existing single-image and video-based pose estimation models but also showcases superior handling of substantial occlusions, achieving fast, robust, accurate, and temporally consistent 3D pose estimates. Code is made publicly available at https://github.com/take2rohit/stride
Figure 1: Effect of occlusions on pose estimation. Image-based 3D pose estimators [3] often struggle with heavy occlusions, as illustrated in this figure. Without temporal context, predictions on highly obscured frames are inconsistent with prior poses, like the erroneous pose in the third column. Notably, even state-of-the-art video approaches [36] fail on prolonged full occlusions spanning multiple frames, as in columns 4 and 5. This highlights yet another critical limitation - models are brittle when deployed outside their training distributions. Without training examples of such long-duration occlusions, models fail to extrapolate reasonable poses. Our work addresses this through test-time training of a human motion prior. By fine-tuning on each new video, we tailor this parametric prior to handling sequence-specific occlusion patterns not observed during training. Given an initial noisy estimate, our approach refines the pose sequence into an accurate, temporally coherent output, as shown in the final row.
* Equal contribution, $\ddagger$ Currently at NASA MSFC IMPACT. Work done while the author was at UCR., $\dagger$ Currently at AWS AI Labs. Work done while the author was at UCR.
@@ 1 Introduction
Accurate 3D pose estimation [48] is an important problem in computer vision with a variety of real-world applications, including but not limited to action recognition [20], virtual and augmented reality [1], and gait recognition [50, 11, 10]. While the performance of 3D pose estimation algorithms has improved rapidly in recent years, the majority of these are image-based [33, 29, 31, 38], estimating the pose from a single image. Consequently, these approaches still face inherent challenges in handling occluded subjects due to the limited visual information contained in individual images. To address these issues, recent efforts have explored video-based pose estimation algorithms [41, 32], leveraging temporal continuity across frames to resolve pose ambiguities from missing visual evidence.
@@
Recent TTT approaches for 3D pose estimation [26, 13, 12] fine-tune models using 2D cues like keypoints from test images. This approach has limitations as the 2D projection of 3D poses is ambiguous, as many 3D configurations can map to the same 2D pose. Also, 2D pose estimators can fail on unseen data distributions [31, 18], so fine-tuning on imperfect and ambiguous 2D poses can lead to incorrect model adaptation and degraded 3D pose predictions.
@@
To overcome the limitations of existing methods, we propose STRIDE (Single-video based TempoRally contInuous Occlusion-Robust 3D Pose Estimation), a novel test-time training framework for 3D pose estimation under occlusion. The key component of our approach is a parametric motion prior that is capable of modeling the dynamics of natural human motions and poses. This motion prior is pre-trained using a BERT-style [9, 51] approach on 3D pose sequences, learning to reconstruct temporally coherent poses when given a series of noisy estimates as input. At test time, given a sequence of noisy 3D poses from any existing pose estimation algorithm, STRIDE leverages this pre-trained prior to produce a clean sequence by fine-tuning it on each new video. We use 3D kinematic losses for motion smoothing via adapting the model to the video-specific motion patterns. By leveraging the motion prior’s inherent knowledge of natural human movement during test-time training, STRIDE avoids ambiguities of 2D pose information faced by existing approaches. An overview of our approach is shown in Fig. 2.
A key advantage of our algorithm is that it can work alongside any off-the-shelf pose estimator to improve temporal consistency, providing model-agnostic pose enhancements. This allows STRIDE to not only surpass image-based pose estimators that lack contextual cues to resolve occlusions, but also outperform video-based methods. Notably, STRIDE can handle situations with up to 100% occlusion of the human body over many consecutive frames. In comparison to existing test-time video based pose estimation method [32, 26], our approach is up to 2$\times$ faster than previous state-of-the-art method [26] and operates without accessing any labeled training data during inference time, making it privacy [34] and storage-friendly. Contributions. In summary, we make the following key contributions:
@@
    We propose a novel test-time training algorithm (STRIDE), for achieving temporally continuous 3D pose estimation under occlusions.
@@
    STRIDE is designed as a model-agnostic framework that leverages a human motion prior model to refine noisy 3D pose sequences from any off-the-shelf estimator into smooth and continuous predictions, highlighting efficiency and generalizability.
@@
    STRIDE achieves state-of-the-art results on multiple challenging benchmarks including OCMotion [15], Occluded Human3.6M and Human3.6M [16], thus demonstrating enhanced occlusion-robustness and temporal consistency. Additionally, STRIDE is computationally amicable, achieving a minimum of 2$\times$ speed-up over existing analogous algorithms.
@@ 2 Related Works
Monocular 3D pose estimation. Monocular 3D pose estimation is a fundamental and challenging problem in computer vision which involves the localisation of 3D spatial pose coordinates from just a single image. Recent deep learning-based methods for the problem have shown impressive performance on challenging academic datasets [3, 48]. [25] introduced the first CNN-based approach to regress 3D joints from a single image. Subsequent works [33, 29] improved upon this by incorporating multi-view constraints and depth information. Recent methods [39, 40] use kinematic and anatomical constraints along with data augmentation to achieve state-of-the-art results on academic datasets. However, these supervised methods often fail under distribution shifts. To address this, [21, 22] proposed self-supervised algorithms for 3D human pose estimation, which perform well on single images but struggle with occlusions and lack temporal continuity in video settings.
@@
Video-based 3D pose estimation. Video-based 3D human pose estimation have shown impressive performance on challenging datasets. [49] directly regresses 3D poses using consistency between 3D joints and 2D keypoints. [30] utilized temporal convolutions for pose estimation in videos, while [2] exploited SMPL pose and shape parameters for fine-tuning HMR in the wild. [44] proposed a mixed spatio-temporal approach alternating between spatial and temporal consistency. HuMoR [32] maintained consistency across frames with weighted regularization using predicted contact probabilities. The state-of-the-art CycleAdapt [26] addresses domain shifts in 3D human mesh reconstruction by cyclically adapting HMRNet [17] and MDNet [26] during test time. Despite the success of the above methods in maintaining temporal consistency, they are extremely slow due to an external optimization step and do not generalise well under distribution shifts. Severe occlusions often degrade the performance of these methods due to missing poses. Our work emphasizes these shortcomings and brings temporal continuity under severe occlusions by leveraging a motion-prior model that seamlessly handles missing poses.
@@
3D pose estimation under occlusion. Handling occlusions poses a significant challenge in both image-based and video-based 3D pose estimation. Approaches like 3DNBF [46] leverage generative models to estimate poses but do not account for any temporal continuity. To alleviate the problem in a video-based setting, [6] introduced data augmentation using occlusion labels with the Cylinder Man Model. Current methods address this by refining 3D poses for temporal consistency. Recent approaches such as GLAMR [41] recover human meshes globally from local motions and perform motion infilling based on visible motions. SmoothNet [42] uses a temporal refinement network to mitigate motion jitters from single image-based pose estimations. While effective for minor occlusions, these methods struggle with heavy occlusions. Also, these algorithms often fail to generalize under domain shifts. To improve on this, our approach adapts a motion prior from noisy 3D pose sequences to predict missing poses and maintain temporal consistency.
Test Time Optimization for 3D pose estimation. A major shortcoming of fully supervised learning is that it can only handle test cases that are similar to the ones seen during the training process. For example, a novel type of occlusion or a human pose during testing can confuse the model and reduce its performance. [43] suggested verifying the estimated 2D poses and additionally ensuring the consistency of the lifted 3D poses by using randomly projected 2D poses to enhance the 3D human pose estimation. Further, [35] introduced the idea of enforcing physical constraints on the estimated human poses to ensure they are physically plausible. Subsequently, [4] proposed combining top-down and bottom-up human pose estimation approaches to take advantage of their strengths and also perform test time optimization using a re-projection loss, and bone length regularizations. Although these methods handle unseen test cases well, they are not designed for handling heavy occlusions and fail to predict poses under complete occlusions. Our work focuses on these gaps by refining and filling in missing 3D poses to maintain temporal consistency.
@@ 3 Method
Figure 3: The presented figure illustrates the pipeline for our temporally continuous pose estimation, STRIDE. Initially, we pre-train a motion prior model, denoted as $\mathcal{M}$, using a diverse set of 3D pose data sourced from various public datasets. The primary objective of this motion prior model is to generate a sequence of poses that exhibit temporal continuity when provided with a sequence of initially noisy poses. Moving into the single video training stage, we acquire a sequence of noisy poses using a 3D pose estimation model, $\mathcal{P}$. The weights of $\mathcal{P}$ are held constant during this phase. Subsequently, we pass this noisy pose sequence through the motion prior model $\mathcal{M}$ and retrain it using various supervised losses, as outlined in Equation 6. The end result of this training process is a model capable of producing temporally continuous 3D poses for that specific video.
We address the problem of extracting temporally continuous 3D pose estimates from a monocular video that may contain heavy occlusions. Given an off-the-shelf monocular 3D pose estimator $\mathcal{P}$ (either image or video-based) that produces temporally inconsistent poses due to occlusions or domain gaps, our goal is to output clean, temporally coherent 3D pose sequences that better match natural human motion dynamics. To achieve this, we propose a two-stage approach, as illustrated in Fig. 3.
@@
    Learning a motion prior: We first pre-train a self-attention-based motion prior model $\mathcal{M}$ on labeled 3D pose datasets in a BERT-style [9, 51]. During pre-training, we synthetically corrupt the 3D joint inputs with noise to simulate occlusions and other errors. $\mathcal{M}$ is then trained to denoise these inputs and reconstruct a sequence of temporally coherent 3D poses. This allows $\mathcal{M}$ to learn strong general priors of natural human motion dynamics.
@@
    Test-time alignment: For a given test video, we obtain noisy per-frame poses using $\mathcal{P}$ [3] and adapt the motion prior model $\mathcal{M}$ in an unsupervised manner to align it to the specific motion exhibited in the video. This adaptation step allows us to obtain optimal pose estimates for the given video.
Section 3.1 outlines the architecture of the motion prior model $\mathcal{M}$. Section 3.2 discusses the masked sequence modelling approach for pre-training $\mathcal{M}$ on synthetically corrupted pose sequences. Lastly, Section 3.3 introduces the self-supervised losses for fine-tuning $\mathcal{M}$ during test time on individual videos.
@@ 3.1 Network Architecture
We base our motion prior model $\mathcal{M}$ on the DSTFormer architecture [51], originally proposed for lifting 2D poses to 3D. Here, we modify and adapt DSTFormer for the sequence-to-sequence task of denoising and smoothing noisy 3D pose sequence inputs. Specifically, the motion prior $\mathcal{M}$ takes in a sequence of 3D body poses represented as $\mathbf{X}\in\mathbb{R}^{T\times J\times 3}$, where $T$ is the number of frames, $J$ is the number of joints, and each pose consists of $J\times 3$ coordinate values. $\mathcal{M}$ denoises the input sequence to produce refined temporally coherent 3D poses $\bar{\mathbf{X}}\in\mathbb{R}^{T\times J\times 3}$. Additional details on the architecture and implementation method is provided in the Section 1 and Section 2 of supplementary material respectively.